# Investigating Primordial Black Hole Accretion through Cosmic Optical Depth

### v1

@@ 5 Discussions
Energy deposition of PBH is dependent on IGM ionization and thermal states, due to observational uncertainties on IGM during reionization, we do not account for PBH energy injection at $z\lesssim 11$, and we ignored contributions from astrophysical sources for convenience. However we note that if both these issues were to be addresses then we expect our constraints to further tighten, thus PBH limits presented here should be considered as conservative. PBH mass growth is expected to continue into even a lower redshift range, particularly the entire cosmic dawn and reionization epoch till around $z\sim 7$. Note that optical depth is an integrated quantity for CMB photon scattering and it does not fully capture the ionization history. Future observations are expected to yield high-precision measurements on $x_{\rm e}(z)$ history [78], which would provide a thorough test on the IGM environmental effect from accelerated PBH radiation. We also note that for $M_{{\rm PBH,ini}}>M_{\odot}$ the stronger accretion radiation also brings more pronounced suppression on the 21-cm absorption signal at early EoR, and may potentially even drive the signal to transition from absorption to emission.

### v2

@@ 5 Discussions
Recently it was pointed out in Ref. [24] that the local ionization and heating can prevent the formation of DM halos surrounding the PBH, potentially leading to reduced accretion rate. Ref. [23] further showed that the formation of ionization fronts can also suppress accretion rate as well as radiation output. These effects can potentially weaken our PBH constraints and we reserve such analysis to future works.
Energy deposition of PBH is dependent on IGM ionization and thermal states, due to observational uncertainties on IGM during reionization, we do not account for PBH energy injection at $z\lesssim 11$, and we ignored contributions from astrophysical sources for convenience. However we note that if both these issues were to be addresses then we expect our constraints to further tighten, thus PBH limits presented here should be considered as conservative. PBH mass growth is expected to continue into even a lower redshift range, particularly the entire cosmic dawn and reionization epoch till around $z\sim 7$. Note that optical depth is an integrated quantity for CMB photon scattering and it does not fully capture the ionization history. In the Planck 2018 results [53], the reionization history was modeled using a phenomenological tanh profile characterized by the reionization midpoint $z_{\rm re}$. Observational uncertainties in $z_{\rm re}$ directly affects the optical depth $\tau$ constraints and consequently our PBH limits. As shown in  [84], low-redshift observations (e.g., galaxy UV luminosity function, Lyman-Alpha forest) can help reconstruct the full reionization history and significantly tighten $z_{\rm re}$ constraints compared to CMB [53]. Incorporating these low-redshift probes can give more robust constraints on PBHs. Future observations are expected to yield high-precision measurements on $x_{\rm e}(z)$ history [85], which would provide a thorough test on the IGM environmental effect from accelerated PBH radiation. We also note that for $M_{{\rm PBH,ini}}>M_{\odot}$ the stronger accretion radiation also brings more pronounced suppression on the 21-cm absorption signal at early EoR, and may potentially even drive the signal to transition from absorption to emission.